OpenStudy (anonymous):

e^x-1-5=5 a.1+ln2+ln5 b.1+ln2-ln5 c.1-ln2+ln5 d.ln2+ln5

OpenStudy (jdoe0001):

\(\large \bf e^{x-1}-5=5 ?\)

geerky42 (geerky42):

Add both sides by 5: \(e^{x-1} = 10\) Take natural log of both sides: \(x-1 = \ln(10) = \ln (5)+\ln(2)\) Add both sides by 1: \(\boxed{x = \ln (5)+\ln(2)+1}\)

OpenStudy (anonymous):

Thank you so much!